OpenStudy (anonymous):

HELPP find the tangents of the curve x^2+ y^2 - 6x + 4y = 0 from (-17, 7). the answers are 2x+3y+13=0 and 34x+129y = 325..

OpenStudy (anonymous):

http://www.weegy.com/

OpenStudy (anonymous):

ha?

OpenStudy (anonymous):

@phi :) this is the last.. very last

OpenStudy (phi):

Here is a picture

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

i messed up again in the cross multipying part.. o men

OpenStudy (phi):

these are wicked. I got \[ \frac{dy}{dx}= \frac{-(x-3)}{(y+2)}\] and the slope equation \[ \frac{-(x-3)}{(y+2)}= \frac{(y-7)}{(x+17)}\] or,after cross multiplying: \[ -(x-3)(x+17)= (y+2)(y-7) \]

OpenStudy (phi):

I see I lost a sign. we work the cross multiply equation into the form \[ x^2+y^2= -14x+5y+65 \] now use the original equation \[ x^2 +y^2= 6x-4y \] \[ 6x-4y = -14x+5y+65 \] solve for y to get \[ y = \frac{20x-65}{9} \]

OpenStudy (phi):

now replace y with that expression in the cross multiply equation As this is very tedious, I used Wolfram to do the hard work. http://www.wolframalpha.com/input/?i=-%28x-3%29%28x%2B17%29%3D+%28y%2B2%29%28y-7%29%2C+y+%3D+%5Cfrac%7B20x-65%7D%7B9%7D